OpenStudy (anonymous):

36^(4c) = √(6)^(8c - 6)

OpenStudy (asnaseer):

hint: write both sides of the equation as 6 to the power of something

OpenStudy (anonymous):

what do you mean?

OpenStudy (asnaseer):

take the 36 on the left hand side for example. how can you write 36 as \(6^a\)? what would a=?

OpenStudy (anonymous):

i mean, i understand that we get sqrt6^(4c)=sqrt(8x-6)

OpenStudy (anonymous):

sqrt6^(8x-6) *

OpenStudy (asnaseer):

no that is not correct

OpenStudy (asnaseer):

\[36=6^a\]what would a=?

OpenStudy (anonymous):

im not sure :/ i thought that at least that way, the sqrt would cancel eachother out and it'd be simpler

OpenStudy (asnaseer):

you don't know how to get 36 as a power of 6?

OpenStudy (asnaseer):

what is \(6\times6=?\)

OpenStudy (anonymous):

ohhhhhhhh, now i get you

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

thanks :)

OpenStudy (asnaseer):

so you know how to solve this now?

OpenStudy (anonymous):

yes, thank you. i wasnt looking for the answer lol. i just needed to remember how to solve it. :)

OpenStudy (asnaseer):

ok - I'm glad I could help. :)